OpenStudy (sloppycanada):

I'm so incredibly lost. Will post below how I think I'm supposed to solve it.

OpenStudy (sloppycanada):

\[\frac{ 9 }{ \sin80 } = \frac{ 5 }{ sinB }\]

OpenStudy (sloppycanada):

http://prntscr.com/9ondop

OpenStudy (sloppycanada):

Sorry, there's the problem.

OpenStudy (sloppycanada):

So from what we learned, I can set the equation up like the above way I had it set up. So don't I just want to divide 9/sin80 by 5?

OpenStudy (sloppycanada):

Am I on the right track or..?

OpenStudy (sloppycanada):

it's flipped, isn't it?

OpenStudy (sloppycanada):

\[\frac{ Sin80 }{ 9 } = \frac{ SinB }{ 5 }\]

OpenStudy (sloppycanada):

\[\sin^-1(\frac{ \sin80(5) }{ 9 }) = measure of ang\]

OpenStudy (sloppycanada):

sorry meassure of angle B?

OpenStudy (dayakar):

change m<A = 80 radians into degrees